OpenStudy (anonymous):

Is there a solution to the following equation on the interval 0 (less than or equal to) theta < (2pi)? sin(2theta)=-(sqrt3)/2

OpenStudy (anonymous):

\[2\theta=n \pi-(-1)^n \pi/6\]

OpenStudy (anonymous):

The angles that generate that sin value are \[4\pi/3\]\[5\pi/3\]

OpenStudy (anonymous):

Ok and which of the two angles do I use in order to solve the equation?

OpenStudy (anonymous):

So the values of x that will work are\[2\pi/3\]\[5\pi/6\]

OpenStudy (anonymous):

Sorry, theta, not x.

OpenStudy (anonymous):